Find the radius of the circle in which a central angle of 60° intercepts an arc of length 37.4 cm (use π = 22/7).

Grade:11

3 Answers

Sathwik reddy
2 years ago
360 angle means 2pi r is length of arc
so for 60 angle the length of arc is pi  r/3 which is equal to 37.4 so
if we do this calculation we get r as some thing which is the answer
 
P S S C ASLESH
2 years ago
Area of sector=lr/2,
wher, l=length of arc and r= radius of circle
Area of sector=(x°/360°)*π
where x°=angle made by the two radii with centre of the circle
i.e; lr/2=(x°/360°)π
      l/2=(x°/360°)πr
     37.4/2=(60/360)*(22/7)r 
     18.7=(1/6)(22/7)r
     18.7*6*7/22=r
      785.4/22=r
     r=35.7cm
Therefore radius of the circle is 35.7cm
 
 
 
 
Please approve it ,if it helps you.
Thank You.
 
Debarati
2 years ago
Let the radius be r
(2*22*r*60)/(7*360) = 37.4 cm
[Since, central angle is 60, we multiply 2*(22/7)*r by 60/360]
=> r = (37.4*7*360)/(2*22*60) = 35.7
So, the radius of the circle is 35.7 cm.